Keith Corporation Balance Sheets December 31 Assets Cash Marketable securitie Accounts receivable Inventories 2019 2018 $1,480 1,770 2,050 2,860 $8,160 $29,470 14,690 $14,780 $22,940 $990 1,210 1,740 2,850 $6,790 $28,150 13,050 $15,100 $21,890 Total current assets Gross fixed assets Less: Accumulated depreciation Net fixed assets Total assets Liabilities and Stockholders Equity Accounts payable Notes payable Accruals $1,640 2,790 200 $4,630 $4,900 S9,530 $10,050 3,360 $13,410 $22,940 $1,510 2,190 300 54,000 $5,070 S9,070 $10,050 2,770 $12,820 $21,890 Total current liabilities Long-term debt otal liabilities Common stock Retained earnings Total stockholders equity Total liabilities and stockholders equity Income Statement Data (2019) Depreciation expense Earnings before interest and taxes (EBIT Interest expense Net profits after taxes Tax rate $1,640 2,710 367 1,851 21%Finding operating and free cash flows Consider the balance sheets and selected data from the income statement of Keith Corporation that follow囲 a. Calculate the firms net operating profit after taxes (NOPAT) for the year ended December 31, 2019. b. Calculate the firms operating cash flow (OCF) for the year ended December 31, 2019. c. Calculate the firms free cash flow (FCF) for the year ended December 31, 2019. d. Interpret, compare and contrast your cash flow estimate in parts (b) and (c)
Answer #1

a) NOPAT = EBIT ( 1- Tax Rate)

NOPAT = 2710*(1-21%)

= 2140.9

b) Operating Cash Flow = Net Income + Non Cash Expenses - Increase in Working Capital

Non Cash Expenses is Depreciation and Amortization

Working Capital = Account Receivable + Inventories - Account Payable

Working Capital in 2019 = 3270

Working Capital in 2018 = 3080

Increase in Working Capital = 190

Operating Cash Flow = 1851 + 1640 - 190 = 3301

c) FCFF = Operating Cash Flow - Capital Expenditure

Capital Expenditure = Gross Fixed Assets 2019 - Gross Fixed Assets 2018 = 1320

FCFF = 3301- 1320 = 1981

d) FCFF gives the more comprehensive outlook of the cash flows of the company. It includes the future capital Expenditures.